# Reward Model Overoptimisation in Iterated RLHF

### v1

@@ 4 Evaluating overoptimisation in Iterated RLHF
4.1 Training setup
Our evaluation setup follows extensive prior works that study overoptimisation in the single iteration RLHF in a controlled and simulated manner coste2024reward ; gao2022scalinglawsrewardmodel . Similarly to coste2024reward we use instructions from the AlpacaFarm dataset dubois2024alpacafarmsimulationframeworkmethods for reward model training and policy optimisation. The preference data $\tilde{\mathcal{D}}_{k}$ collected at each iteration contains preferences for a subset of $1000$ instructions in the preference split of AlpacaFarm. Preference labels $p_{i}$ are simulated with the 7 billion parameter Reward-Model-AlpacaFarm-Human dubois2024alpacafarmsimulationframeworkmethods , which is also used by coste2024reward . It is worth noting again the significant difference in parameter sizes between the proxy reward models and the gold reward model, justifying the use of the gold reward model as a proxy for human labellers. Similarly to coste2024reward , to obtain $\pi^{sft}$, we performed supervised fine-tuning on the pythia-$410$m model [^1] https://huggingface.co/EleutherAI/pythia-410m on the AlpacaFarm SFT split. We chose pythia-$410$m as it achieves an appropriate balance between computational cost and experimental rigour for our investigation. The authors in gao2022scalinglawsrewardmodel also found that policy size did not affect the shape of the overoptimisation curve in their setting, further justifying this choice of policy. We initialise proxy reward models $\tilde{R}_{k}$ from the huggingface checkpoint tlc4418/pythia_70m_sft provided by coste2024reward  [^2] https://huggingface.co/tlc4418/pythia_70m_sft, as well as the larger pythia-$160$m, with a randomly initialised prediction head coste2024reward . We train reward models for $5$ epochs with a learning rate of $1\times 10^{-5}$ coste2024reward . For policy optimisation, we perform $6000$ steps of PPO on the unlabelled split of AlpacaFarm. The learning rate is set to $1\times 10^{-6}$ and a constant KL penalty of $1\times 10^{-4}$ is used. The full specifications of the hyperparameters for reward model training and policy optimisation, and the prompt format are given in Appendix C.
@@ 4.1 Training setup
4.2 Measuring reward model overoptimisation with the Maximum Mean Discrepancy
The standard methodology for investigating reward model overoptimisation is to compare the achieved mean reward on the proxy and gold reward functions on a hold-out set coste2024reward ; moskovitz2023confrontingrewardmodeloveroptimization ; gao2022scalinglawsrewardmodel . However, this analysis does not capture discrepancies in the long-tail, i.e., in the high-reward regime, that have a larger impact on the policy optimisation. In this work we propose to compare two reward models based on their distributions of rewards. To this end, we evaluate the policy and reward models on the $2000$ unseen instructions contained in the validation split of AlpacaFarm at every $300$ steps during policy optimisation.
Our approach to measuring differences between reward functions consists of two steps, the first of which is a standardisation that ensures reward functions that lead to the same ordering of policies when optimised are treated as equal (see Apendix B). In the second step, we use the maximum mean discrepancy (MMD) gretton2012MMD to measure the discrepancy between the two reward functions. In particular, we utilise this method to compare the proxy reward models trained at each iteration with the gold-reward model $R^{\star}$. For full details and a justification of the validity of this method we refer the reader to Appendix A.
@@ 5.1 Iterated RLHF can close the gap between proxy and gold reward functions
Furthermore, the KL-reward Pareto front advances with each iteration, although improvements diminish as the distance between proxy and gold reward curves shrinks in later iterations. Despite these diminishing returns, there remains scope to better align gold and proxy reward functions. Comparing reward distributions across iterations further reveals that, after the policy closely approximates the output distribution in $\mathcal{D}_{k}$, the MMD increases again in the high-KL regime for all iterations, especially rapidly in the initial iteration (see Figure 4). Additionally, the rate at which the proxy-gold reward gap closes varies considerably among methods (see Appendix E.1), highlighting the importance of investigating design choices described in Section 3.
Figure 4: Progression of proxy and gold reward alignment across RLHF iterations with policy reinitialization from $\pi^{\text{sft}}$ and concatenated preference data. Mean gold and proxy scores (with standard error across seeds) indicate a narrowing gap across iterations, demonstrating improved robustness of the proxy reward model and alignment with the gold reward. However, diminishing returns are visible in later iterations. Maximum Mean Discrepancy (MMD) between proxy and gold reward distributions on the hold-out set shows initial convergence, but reveals a consistent rise at higher KL values, especially in early iterations, highlighting persistent distributional shifts and misalignments not captured by mean scores alone.
@@ 5.2 Combining preference data
Scaling reward model training data is most effective. We first focus on comparing methods for combining preference datasets. To isolate the effects of varying the combination strategy, we fix the policy initialization to From SFT (Figure 3.i) and reward models are combined using the Take last (Figure 3.d) approach. As shown in Figure 5(a), all methods demonstrate significant improvements over a single iteration, particularly in preventing performance collapse at higher KL divergences.
The Concatenate (Figure 3.b) strategy achieves consistently higher Gold Scores, especially in the KL range of 50-200 nats, where it maintains a performance advantage over other methods. While Take last (Figure 3.a) and Sample (Figure 3.c) approaches show similar trends and substantial improvements over iteration 1, they do not quite match the performance of full data concatenation. This result is coherent with the finding that increasing training dataset size reduces reward model overoptimisation gao2022scalinglawsrewardmodel , explaining why the sampling strategy is outperformed by concatenating all datasets. A critical observation is that beyond $\mathrm{KL}\approx 200$, the baseline iteration 1 experiences severe performance degradation due to overoptimisation, dropping to negative gold scores. In contrast, all iterative approaches maintain positive performance even at high KL values, demonstrating their effectiveness in mitigating overoptimisation. This ranking of methods is not only observed in the final iteration, but is already exhibited as early as the second iteration as shown in Figure 2 and in Appendix E.2.
Ensuring full coverage of the prompts when sampling matters less. While the sampling strategy slightly outperformed taking only the newest preference dataset, it did not achieve the same level of performance as concatenating all data. Here we take a closer look at the sampling strategy. In Figure 5(b) standard sampling with potential prompt repetition (Sample) and sampling where each prompt appears exactly once Sample Exclusive. While Sample Exclusive shows a slight advantage in the mid-KL regime (100-250 nats), the differences are modest, suggesting that prompt repetition plays a minor role in model performance and mitigating overoptimisation. These observations also hold throughout the earlier iterations (see Appendix E.2), highlighting that while data combination strategies are effective at preventing overoptimisation, the computational cost of maintaining and training on growing datasets remains a necessity as more efficient methods are unable to achieve the same performance as Concatenate. This motivates our exploration of alternative approaches focusing on the combination of reward models in the parameter space, which could potentially offer similar benefits with reduced computational overhead.
@@
Figure 5: Iterated RLHF benefits most from scaling reward model training data. (a) Concatenating all preference data across iterations consistently outperforms sampling or using only recent data, particularly at mid KL values (50–200 nats), and best mitigates overoptimisation. (b) Sampling with or without prompt repetition yields similar results, suggesting prompt coverage has limited impact. (c) Parameter-space reward model combinations like ensembling and weight averaging offer efficiency gains but do not surpass the simpler Take last approach. Concatenation remains most effective, motivating efficient alternatives.
@@ 5.3 Combining reward models
Figure 6: Impact of reward model size and design choices on gold score across iterations. Larger reward models (160M, solid lines) consistently outperform smaller ones (70M, dashed), with the biggest gains seen in Ensemble and Worst-Case Optimisation—suggesting that these strategies benefit most from increased reward model capacity. While From SFT remains the most stable across scales, LITI shows steady improvement, especially with the larger model.
No free lunch by merging reward models. The method of concatenating all preference data, previously shown to be most effective, represents the level of performance we aim to achieve. As shown in Figure 5(c), all approaches demonstrate similar patterns of improvement in the early KL regions (up to $\approx 50$ nats), reaching comparable performance levels. The Weight Average (Figure 3.f) and Ensemble (Figure 3.e) methods show consistent performance throughout, while maintaining computational efficiency compared to the data concatenation approach. However, using a reward model ensemble scales inference time and memory requirements during policy optimisation. We observe the mean objective does not yield significant performance gains over the Take Last approach relying on a single reward model (consistent with the findings for the single iteration setting in coste2024reward ). While weight averaged reward models have been shown to outperform ensembles ramé2024warm , we do not observe significant differences except on the mid-KL regime. In contrast to coste2024reward ; ramé2024warm , we ensemble and weight average models that have been trained on data with significantly different joint distribution over pairs $(x,y)$. Regardless, both methods still provide significant improvements when comparing the fourth and first iterations. The various reward model combination methods in RLHF perform similarly, suggesting computational efficiency should drive selection.
@@ 5.4 Policy initialisation
Overoptimised policies are hard to recover from. While From SFT is reset at the end of each iteration, LITI and Take last have to recover form the initial overoptimisation, as shown in Figure 7(b). The behaviour in earlier iterations reveals the consistent performance improvements attained with LITI. On the other hand, Take last is unable to recover after overoptimising again in the second iteration, despite the counterpart – sampling preference data but initialising From SFT – improving with each iteration. Due to entropy decreasing caused by to the prolonged optimisation when using the Take last initialisation, the mean gold reward does not exceed zero in the third and fourth iterations. Despite LITI improving on average across multiple seeds, we observe that linear interpolation is also unable to recover strongly overoptimised seeds (see Appendix E.4). Thus, while From SFT is most robust, it is also limited by the repeated initialisation from $\pi^{sft}$.
Policy interpolation works better with larger reward models. We hypothesise that LITI could achieve similar or higher gold scores than From SFT after more iterations. Supporting this, our experiments with a larger reward model show that LITI benefits substantially from increased reward model capacity (see Figure 6). This improvement likely stems both from better-calibrated gradients that support recovery, and from the fact that larger reward models tend to overoptimise less aggressively (gao2022scalinglawsrewardmodel, ), resulting in safer intermediate policies and more stable interpolation paths. These findings highlight the importance of early stopping and reward model design when using policy initialisation methods other than From SFT, and suggest that LITI may become increasingly competitive as reward model expressiveness scales.
@@
Figure 7: Effect of policy initialisation on overoptimisation and recovery across iterations. At iteration 4, initialising From SFT is most robust, avoiding overoptimisation through repeated resets and data aggregation. In contrast, LITI and Take last start from high KL due to cumulative divergence. When Plotting reward against timesteps and across iterations we observe that overoptimised policies are hard to recover: Take last degrades further, while LITI shows steady improvement, though neither fully closes the gap to From SFT.

### v2

@@ 4 Evaluating overoptimisation in Iterated RLHF
Training setup. Our evaluation setup follows extensive prior works that study overoptimisation in the single iteration RLHF in a controlled and simulated manner (Coste et al., 2024; Gao et al., 2023). Similarly to Coste et al. (2024) we use instructions from the AlpacaFarm dataset (Dubois et al., 2023) for reward model training and policy optimisation. The preference data $\tilde{\mathcal{D}}_{k}$ collected at each iteration contains preferences for a subset of $1000$ instructions in the preference split of AlpacaFarm. Preference labels $p_{i}$ are simulated with the 7 billion parameter Reward-Model-AlpacaFarm-Human (Dubois et al., 2023), which is also used by Coste et al. (2024). It is worth noting again the significant difference in parameter sizes between the proxy reward models and the gold reward model, justifying the use of the gold reward model as a proxy for human labellers. Similarly to Coste et al. (2024), to obtain $\pi^{sft}$, we performed supervised fine-tuning on the pythia-410410m model (Biderman et al., 2023) on the AlpacaFarm SFT split. We chose pythia-410410m as it achieves an appropriate balance between computational cost and experimental rigour for our investigation. Gao et al. (2023) also found that policy size did not affect the shape of the overoptimisation curve in their setting, further justifying this choice of policy. We initialise proxy reward models $\tilde{R}_{k}$ from the HuggingFace checkpoint pythia_70m_sft provided by Coste et al. (2024), as well as the larger pythia-160160m, with a randomly initialised prediction head (Coste et al., 2024). We train reward models for $5$ epochs with a learning rate of $1\times 10^{-5}$ (Coste et al., 2024). For policy optimisation, we perform $6000$ steps of PPO on the unlabelled split of AlpacaFarm. The learning rate is set to $1\times 10^{-6}$ and a constant KL penalty of $1\times 10^{-4}$ is used. The full specifications of the hyperparameters for reward model training and policy optimisation, and the prompt format are given in Appendix C.
@@
Measuring overoptimisation with the Maximum Mean Discrepancy. The standard methodology for investigating reward model overoptimisation is to compare mean rewards on proxy vs. gold reward functions over a hold-out set (Coste et al., 2024; Moskovitz et al., 2024; Gao et al., 2023). This overlooks discrepancies in the high-reward tail, which more strongly influence policy optimisation. We instead compare reward models by their distributions of rewards, evaluating on the $2000$ unseen instructions contained in the validation split of AlpacaFarm at every $300$ steps during policy optimisation.
Our approach to measuring differences between reward functions consists of two steps, the first of which is a standardisation that ensures reward functions that lead to the same ordering of policies when optimised are treated as equal (see Appendix B.1). In the second step, we use the maximum mean discrepancy (MMD) (Gretton et al., 2012) to measure the discrepancy between the two reward functions. In particular, we utilise this method to compare the proxy reward models trained at each iteration with the gold-reward model $R^{\star}$. For full details and a justification of the validity of this method we refer the reader to Appendix B.
@@ 5.1 Iterated RLHF can close the gap between proxy and gold reward functions
Furthermore, the KL-reward Pareto front advances with each iteration, although improvements plateau as the distance between proxy and gold reward curves shrinks in later iterations. These performance plateaus appear to result from a combination of interacting factors rather than simple diminishing returns. First, the proxy reward model progressively converges toward the gold reward model on the distribution induced by policy optimisation, which limits the scope for further improvement. Second, policy entropy tends to decline across iterations, particularly when initialisation methods other than From SFT are used. Third, data saturation may occur once additional preference data provides little novel information. However, there remains scope to better align gold and proxy reward functions. Comparing reward distributions across iterations further reveals that, after the policy closely approximates the output distribution in $\mathcal{D}_{k}$, the MMD increases again in the high-KL regime for all iterations, especially rapidly in the initial iteration (see Figure 4). We hypothesize that the non-monotonic relationship between MMD and KL reflects a dynamic interplay between alignment and exploitation during training. Early on, RL against the proxy RM improves alignment with held-out samples from initialisation, reducing MMD as the proxy’s predictions grow closer to the gold RM. Later, as the policy distribution shifts and begins exploiting proxy-specific quirks (increasing KL), outputs diverge from true human preferences, driving MMD back up. Additionally, the rate at which the proxy-gold reward gap closes varies considerably among methods (see Appendix D.1), highlighting the importance of investigating design choices described in Section 3.
Figure 4: Progression of proxy–gold alignment across RLHF iterations with $\pi^{\text{sft}}$ reinitialisation and concatenated data. Mean scores show narrowing gaps and improved robustness, though with diminishing returns. MMD reveals early convergence but rising divergence at higher KL, highlighting distributional shifts not observed in mean scores.
@@ 5.2 Combining preference data
Scaling reward model training data is most effective. We first focus on comparing methods for combining preference datasets. To isolate the effects of varying the combination strategy, we fix the policy initialization to From SFT and reward models are combined using the Take last approach. As shown in Figure 5(a), all methods demonstrate significant improvements over a single iteration, particularly in preventing performance collapse at higher KL divergences.
The Concatenate strategy achieves consistently higher gold scores, especially in the KL range of 50-200 nats. While Take last and Sample approaches show similar trends and substantial improvements over iteration 1, they do not quite match the performance of full data concatenation. This result is coherent with the finding that increasing training dataset size reduces reward model overoptimisation (Gao et al., 2023), explaining why the sampling strategy is outperformed by concatenating all datasets. A critical observation is that beyond $\mathrm{KL}\approx 200$, the baseline iteration 1 experiences severe performance degradation due to overoptimisation, dropping to negative gold scores. In contrast, all iterative approaches maintain positive performance even at high KL values, demonstrating their effectiveness in mitigating overoptimisation. This ranking of methods is not only observed in the final iteration, but is already exhibited as early as the second iteration as shown in Figure 2 and in Appendix D.2.
Ensuring full coverage of the prompts when sampling matters less. While the sampling strategy slightly outperformed taking only the newest preference dataset, it did not achieve the same level of performance as concatenating all data. Here we take a closer look at the sampling strategy. In Figure 5(b) standard sampling with potential prompt repetition (Sample) and sampling where each prompt appears exactly once (Sample Exclusive). The differences are minor, suggesting that prompt repetition has a limited impact on performance or overoptimisation. This pattern also holds in earlier iterations (Appendix D.2), highlighting that while data combination strategies are effective at preventing overoptimisation, the computational cost of maintaining and training on growing datasets remains, as more efficient methods are unable to achieve the same performance as Concatenate. This motivates exploring reward-model combination in parameter space to achieve similar gains with less overhead.
@@
Figure 5: Iterated RLHF benefits most from scaling reward model training data. (a) Concatenating all preference data across iterations best mitigates overoptimisation, especially at mid KL (50–200). (b) Sampling, with or without prompt repetition, performs similarly, implying limited impact of prompt coverage. (c) Parameter-space methods (ensembles, averaging) lead to efficiency gains but fall short of the simpler Take last with data aggregation.
@@ 5.3 Combining reward models
No free lunch by merging reward models. Concatenating all preference data, previously the most effective method, serves as our performance baseline. As shown in Figure 5(c), all approaches improve similarly in early KL regions (up to $\approx 50$ nats), reaching comparable performance. Weight Average and Ensemble maintain strong, efficient performance, though ensembles increase inference time and memory use. The mean objective offers no clear gains over the Take Last approach with a single reward model, consistent with Coste et al. (2024). While weight averaging has been reported to outperform ensembles (Ramé et al., 2024b), we only observe differences in the mid-KL regime. In contrast to prior work (Coste et al., 2024; Ramé et al., 2024b), we combine models trained on data with significantly different joint distribution over pairs $(x,y)$. Regardless, both methods still provide significant improvements when comparing the fourth and first iterations. The various reward model combination methods in RLHF perform similarly, suggesting computational efficiency should drive selection.
Figure 6: Impact of reward model size on gold score. Larger models (160M, solid) outperform smaller ones (70M, dashed), with the biggest gains in Ensemble and Worst-Case Optimisation. From SFT stays stable, while LITI steadily improves with scale.
@@ 5.4 Policy initialisation
Overoptimised policies are hard to recover from. While From SFT is reset at the end of each iteration, LITI and Take last have to recover form the initial overoptimisation, as shown in Figure 7(b). The behaviour in earlier iterations reveals the consistent performance improvements attained with LITI. On the other hand, Take last is unable to recover after overoptimising again in the second iteration, despite the counterpart, sampling preference data but initialising From SFT, improving with each iteration. Due to entropy decreasing caused by to the prolonged optimisation when using the Take last initialisation, the mean gold reward does not exceed zero in the third and fourth iterations. Despite LITI improving on average across multiple seeds, we observe that linear interpolation is also unable to recover strongly overoptimised seeds (see Appendix D.4). Thus, while From SFT is most robust, it is also limited by the repeated initialisation from $\pi^{sft}$.
Policy interpolation works better with larger reward models. We hypothesise that LITI could achieve similar or higher gold scores than From SFT after more iterations. Supporting this, our experiments with a larger reward model show that LITI benefits substantially from increased reward model capacity (see Figure 6). This improvement likely stems both from better-calibrated gradients that support recovery, and from the fact that larger reward models tend to overoptimise less aggressively (Gao et al., 2023), resulting in safer intermediate policies and more stable interpolation paths. These findings highlight the importance of early stopping and reward model design when using policy initialisation methods other than From SFT, and suggest that LITI may become increasingly competitive as reward model expressiveness scales.
@@
Figure 7: Effect of policy initialisation on overoptimisation and recovery across iterations. From SFT is most robust, avoiding divergence via resets and data aggregation. LITI and Take last start from high KL due to cumulative divergence. Overoptimised policies are hard to recover: Take last worsens over time, while LITI improves but does not reach From SFT.
6 Limitations
Our study focuses on controlled settings, using modestly sized policy models (Pythia-410M) and reward models (70M, 160M) on the AlpacaFarm benchmark with a static “gold” reward model to simulate human feedback. However, this setup, consistent with prior work (Coste et al., 2024; Ramé et al., 2024b; Zhu et al., 2024) (also in terms of model size), enables systematic investigation of iterative RLHF while ensuring results remain interpretable and comparable. Moreover, scaling laws suggest policy size is not the main driver of overoptimisation and that scale effects are smooth (Gao et al., 2023), indicating that our findings and design choices are likely to transfer to larger models, even if the degree or speed of overoptimisation differs. Using a single dataset (AlpacaFarm) enabled controlled ablations but may not capture the diversity of real-world tasks. We also note that a static “gold” reward model, abstracts away the noisy and evolving nature of human preferences. However, this is standard practice in the field (Coste et al., 2024; Gao et al., 2023) to ensure reproducibility and mitigate cost of human labelling. We ran four iterations, enough to observe plateaus and overoptimisation trends, but further scaling was prohibitive given the computational resources of our institution. Nonetheless, our work offers key insights and guidance for designing more robust iterative RLHF pipelines and lays groundwork for future research on larger scales and real-world settings.
@@ Acknowledgements
Appendix A A theoretical framework: Iterated RLHF and Performative Prediction
A.1 Overview
We note that the framework of performative prediction (Perdomo et al., 2020) can be applied to our setting. In fact, when performing iterated RLHF, we are simulating performative prediction or more specifically a version of strategic classification. We have that a reward model $R_{\phi}$ induces a potentially different distribution $\mathcal{D(\phi)}$ over instances $(x,y)$ where continuations $y$ are obtained from the policy $\pi_{\theta}$ optimised for $R_{\phi}$, which yields that a reward model $R_{\phi_{PO}}$ is performatively optimal if $\phi_{PO}=\underset{\phi}{\arg\min}\mathbb{E}_{(x,y)\sim\mathcal{D(\phi)}}\ell((x,y,\phi))$. Furthermore, a model $R_{\phi_{PS}}$ is defined as performatively stable if $\phi_{PS}=\underset{\phi}{\arg\min}\mathbb{E}_{(x,y)\sim\mathcal{D}(\phi_{PS})}\ell((x,y,\phi)).$ Intuitively, retraining a performatively stable reward model after optimising against it will yield the same reward model. As such the reward model would not be over-optimised and still perform optimally on its induced distribution. In Theorem 3.5 Perdomo et al. (2020) provide 3 conditions under which the reward model obtained from repeated iterations of RLHF converges to a unique performatively stable reward model at a linear rate. We require the loss to be $\beta$-jointly smooth and $\gamma$-strongly convex, and the map $\mathcal{D}(\cdot)$ from reward model parameters to the distribution of prompt continuation pairs to be $\epsilon$-sensitive. Since as part of the map $\mathcal{D}(\cdot)$ the policy is optimised with PPO, where small changes in the reward model can lead to significant changes in the optimal policy, this mapping is generally not $\epsilon$-sensitive. As a consequence, linear convergence is not guaranteed. Note, that we may still aim for close to linear convergence by making adjustments to satisfy the stated conditions.
In the following subsections we expand on the overview above and present a concise theoretical account of iterated RLHF. We use $\phi$ to denote reward‑model parameters and $\theta$ to denote policy parameters. Our presentation casts iterated RLHF as a performative prediction problem, and then derives sufficient conditions for convergence as well as a set of practical propositions that explain common empirical mitigations (data aggregation, reward model ensembles, and policy resetting).
A.2 Setup
Let $\pi_{\theta}$ be a stochastic policy parameterised by $\theta\in\Theta$. Let $R_{\phi}$ be a learned reward model parameterised by $\phi\in\Phi$. We denote by $\Pi(R_{\phi})\mapsto\pi_{\theta(\phi)}$ the policy optimisation operator that (approximately) returns a policy optimised with respect to $R_{\phi}$. Running $\pi_{\theta(\phi)}$ in the environment (or simulator) induces a distribution over prompts and model responses; we write $D(\phi)$ for the resulting distribution over observed preference pairs or (input, response) tuples $(x,y)$.
A reward model is trained by empirical risk minimization on data sampled from the distribution induced by the current reward model through policy optimisation. Concretely, given a loss function $\ell(\phi;(x,y))$ (for example cross‑entropy or a surrogate for pairwise preference loss), the standard iterated update considered throughout this work can be defined as follows:
$\phi_{t+1}=\arg\min_{\phi}\;\mathbb{E}_{(x,y)\sim D(\phi_{t})}\big[\ell(\phi;(x,y))\big].$ (3)
This framing matches the performative prediction viewpoint: the object being learned (the reward model) affects the data distribution through the downstream policy it induces.
A.3 Performative stability
Definition A.1 (Performative stability).
A reward model parameter $\phi^{*}$ is called performatively stable if it is a fixed point of the update equation 3, i.e.
$\phi^{*}=\arg\min_{\phi}\;\mathbb{E}_{(x,y)\sim D(\phi^{*})}\big[\ell(\phi;(x,y))\big].$
At a performatively stable point, retraining the reward model on data produced by the policy it induces produces no change. Iterated RLHF can therefore be interpreted as an algorithmic attempt to reach such a fixed point.
A.4 Convergence guarantees
Theorem A.2 (Convergence to a performatively stable point).
Suppose the per‑example loss $\ell(\phi;(x,y))$ is $\alpha$-strongly convex and $\beta$-smooth in $\phi$ for every data point $(x,y)$. Suppose further that the mapping $\phi\mapsto D(\phi)$ is $L$-Lipschitz in total variation distance. If $\frac{L\beta}{\alpha}<1,$ then the map defined by the update equation 3 is a contraction and the iterates $\{\phi_{t}\}_{t\geq 0}$ converge linearly to a unique performatively stable point $\phi^{*}$.
Proof sketch. For each fixed data distribution, strong convexity and smoothness imply the population risk admits a unique minimizer, and the $\arg\min$ mapping is Lipschitz with constant at most $\beta/\alpha$. Composing this with the $L$-Lipschitz dependence of $D(\phi)$ on $\phi$ yields an overall contraction constant bounded by $L\beta/\alpha$. If this constant is strictly less than one, Banach’s fixed point theorem guarantees a unique fixed point and geometric convergence of iterates. This is an application of the performative prediction contraction framework (Perdomo et al., 2020).
Discussion. The theorem isolates two failure modes in practice: (i) the loss used to train reward models is rarely globally strongly convex in modern neural parameterisations, and (ii) modern policy optimisers (PPO, SAC, etc.) can induce highly non‑Lipschitz changes in the data distribution, i.e., small changes to $\phi$ may yield large shifts in $\pi_{\theta(\phi)}$ and hence in $D(\phi)$. Consequently, the sufficient conditions above are not satisfied in general RLHF pipelines, but they nevertheless clarify why certain regularisers and protections (e.g., constraining policy updates, aggregating data) promote stable behaviour.
A.5 Preference data aggregation
Proposition A.3 (Data aggregation reduces estimation error).
Let the reward model be trained by empirical risk minimization on a dataset $\mathcal{S}$ of size $N$. Under standard i.i.d. concentration bounds, the expected generalization error of the empirical minimizer scales as $O(1/\sqrt{N})$. If datasets collected across iterations $\mathcal{S}_{1},\dots,\mathcal{S}_{T}$ are concatenated to form $\mathcal{S}_{\mathrm{tot}}$ with total size $N_{\mathrm{tot}}=\sum_{t=1}^{T}N_{t}$, the estimation error correspondingly decreases as $O(1/\sqrt{N_{\mathrm{tot}}})$.
Proof sketch. This follows from Hoeffding‑type concentration or uniform convergence arguments: more samples tighten empirical estimates of the population risk and hence reduce the gap between empirical and population minima.
Corollary A.4.
Training on aggregated data approximates training on the mixture distribution $D_{\mathrm{mix}}=\frac{1}{T}\sum_{t=1}^{T}D(\phi_{t})$, reducing variance and decreasing sensitivity to idiosyncrasies of any single iteration.
Discussion. Aggregation stabilizes training in two ways: it increases effective sample size (reducing estimation noise) and smooths the effective data generating process, which can reduce the Lipschitz constant of $\phi\mapsto D(\phi)$ empirically.
A.6 Reward‑model ensembles and transfer
Proposition A.5 (Averaging reduces squared error).
Let $R_{\phi_{i}}=R^{*}+\delta_{i}$ be $K$ proxy reward models with additive errors $\delta_{i}$. Let define the ensemble reward as: $R_{\mathrm{ens}}=\frac{1}{K}\sum_{i=1}^{K}R_{\phi_{i}}=R^{*}+\frac{1}{K}\sum_{i=1}^{K}\delta_{i}.$ Then
$|R_{\mathrm{ens}}-R^{*}|_{2}^{2}=\Big|\frac{1}{K}\sum_{i=1}^{K}\delta_{i}\Big|_{2}^{2}\leq\frac{1}{K}\sum_{i=1}^{K}|\delta_{i}|_{2}^{2}.$
Proof sketch. This is a direct consequence of Jensen’s inequality / the variance reduction property of averaging.
Discussion. When errors $\delta_{i}$ are approximately zero‑mean and weakly correlated, ensembles can substantially reduce the magnitude of systematic errors that policies can exploit. Worst‑case ensemble strategies (e.g., conservative lower‑bound ensembles) further limit reward overestimation.
A.7 Policy initialization and reset strategies
Let $\pi_{\theta_{0}}$ denote a base supervised fine‑tuned (SFT) policy. Let us define the Kullback–Leibler divergence between two policies by $D_{\mathrm{KL}}(\pi_{\theta}\|\pi_{\theta_{0}})$.
Proposition A.6 (Resetting bounds policy drift).
If at every iteration the policy optimisation is initialized from the base policy $\pi_{\theta_{0}}$ (i.e. we re‑start optimisation from $\theta_{0}$), then the accumulated divergence from the base policy over iterations is bounded by the per‑iteration optimisation step sizes. In contrast, warm‑starting from the previous iterate $\theta_{t-1}$ can lead to cumulative drift: divergences can add across iterations and become large.
Discussion. Resetting is an effective empirical safeguard against runaway behaviour and can improve reproducibility at the cost of reduced per‑iteration adaptivity.
A.8 Overoptimisation (error‑to‑gap) bound
Proposition A.7 (Error–to–gap bound).
Suppose the reward model approximation error is uniformly bounded: for all outputs $y$, $|R_{\phi}(y)-R^{*}(y)|\leq\varepsilon.$
Then the suboptimality gap in the maximized rewards satisfies
$\max_{y}R^{*}(y)-\varepsilon\leq\max_{y}R_{\phi}(y)\leq\max_{y}R^{*}(y)+\varepsilon.$
Discussion. Bounding the sup‑norm error of the reward model controls the extent to which an optimiser can overestimate the true reward. The preceding propositions (aggregation and ensembling) are practical mechanisms for reducing $\varepsilon$ and hence for limiting overoptimisation.
A.9 Concluding Remarks
Framing iterated RLHF as a performative prediction problem clarifies both desirable algorithmic choices and structural failure modes. Under favourable convexity and Lipschitz conditions one recovers a contraction argument guaranteeing convergence to a unique performatively stable reward model. In realistic RLHF pipelines these conditions fail, but the theory explains why mitigation strategies—data aggregation, reward model ensembles, and policy resets—improve stability: they reduce estimation variance, shrink reward‑model error, and bound policy drift. Together these tools help iterated RLHF approximate performatively stable equilibria in practice, even when exact theoretical conditions are not met.